# Detection of evolutionary shifts in variance under an Ornstein–Uhlenbeck model

### v2

@@ 2.2 Evolutionary shifts in optimal value and variance
Let $\gamma_{b}=\Delta\sigma_{b}^{2}$, $\mathbf{V}_{i,j}=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}(1-e^{-2\alpha t_{a}})$, $\mathbf{U}_{i,j}=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}$, $q_{b}=\frac{e^{-2\alpha T}}{2\alpha}\left(1-e^{2\alpha t_{b}}\right)$; $\gamma_{b}$ is the shift in variance on branch $b$, $\mathbf{V}$ is the phylogenetic covariance matrix when $\sigma^{2}=1$ (no shift in variance) with fixed root, $\mathbf{U}$ is the phylogenetic covariance matrix when $\sigma^{2}=1$ (no shift in variance) with stationary distributed root. $\mathbf{U}$ and $\mathbf{V}$ only depend on the phylogeny. The covariance between species $i$ and $j$ can be expressed as the following equation.
$\Sigma_{i,j}=\begin{cases}\mathbf{V}_{i,j}\sigma_{0}^{2}+\sum_{b}\gamma_{b}X_{ib}X_{jb}\mathbf{V}_{i,j}+\sum_{b}\gamma_{b}q_{b}X_{ib}X_{jb}&\text{OU model with fixed root}\\ \mathbf{U}_{i,j}\sigma_{0}^{2}+\sum_{b}\gamma_{b}X_{ib}X_{jb}\mathbf{V}_{i,j}+\sum_{b}\gamma_{b}q_{b}X_{ib}X_{jb}&\text{OU model with random root}\end{cases}$
It consists of 3 terms. The first term is the original covariance without any shift. The last 2 terms show the influence of shifts in variances. In addition to the shift sizes $\gamma_{b}$, the second term is influenced by the phylogentic structure ($\mathbf{V}_{i,j}$), higher original covariance between two species leads to larger change in covariance for a fixed shift size (Fig. 2); the third term is influenced by the start time of branch b ($\mathbf{q}_{b}$), earlier shifts lead to larger change in covariance between two species (Fig. 2).
Figure 1: Higher original covariance between two species leads to larger change in covariance for the same shift magnitude: larger change for the shift in the left figure
Figure 2: Earlier shift leads to larger change in covariance for a given species: larger change for the shift in the left figure
@@
$=\sigma_{0}^{2}\mathbf{R}+\left(\mathbf{X}\operatorname{diag}(\gamma)\mathbf{X^{T}}\right)\odot\mathbf{V}+\mathbf{X}\operatorname{diag}(\gamma\odot q)\mathbf{X^{T}}$ (3)
Where $\odot$ represents elementwise multiplication of matrices and vectors. The trait values at tips can be written as:
@@ 3 Methods
$=-\frac{1}{2}(\mathbf{Y}-\beta_{0}-\mathbf{X}\mathbf{\beta})^{T}\left(\sigma_{0}^{2}\mathbf{R}+\left(\mathbf{X}\operatorname{diag}(\gamma)\mathbf{X^{T}}\right)\odot\mathbf{V}+\mathbf{X}\operatorname{diag}(\gamma\odot q)\mathbf{X^{T}}\right)^{-1}(\mathbf{Y}-\beta_{0}-\mathbf{X}\mathbf{\beta})$
$\qquad-\frac{1}{2}\log{\det{\left(\sigma_{0}^{2}\mathbf{R}+\left(\mathbf{X}\operatorname{diag}(\gamma)\mathbf{X^{T}}\right)\odot\mathbf{V}+\mathbf{X}\operatorname{diag}(\gamma\odot q)\mathbf{X^{T}}\right)}}$
To select the shifts in optimal values ($\beta_{i}\neq 0$) and the shifts in variances ($\gamma_{i}\neq 0$), we use an L1 penalty in the loss function to conduct the feature selection as in LASSO. Therefore, the loss function to be optimized is given by:
@@
$\qquad+\lambda_{1}\|\beta\|_{1}+\lambda_{2}\|\gamma\|_{1}$ (4)
@@ 3.1 Optimization
In this paper, we do not estimate the parameter $\alpha$ jointly with other parameters, but instead treat it as fixed using an ad-hoc estimate obtained from the null model fitted by the phylolm R package. Therefore, in this section, we treat it as fixed during the model fitting process. In the simulations, we demonstrate that the estimation of $\alpha$ does not significantly affect the detection results. Therefore, we omit the estimation of $\alpha$ in the ShiVa procedure. In cases where $\alpha$ is unknown, we use an ad-hoc estimate of $\alpha$ provided by fitting a null model (a model without shifts) using the phylolm package. This estimated $\alpha$ is then fixed when running ShiVa. The primary objective here is to find the parameters $\mathbf{\beta}$, $\mathbf{\gamma}$, and $\sigma_{0}^{2}$ that minimize the loss function described in Equation 3. When $\mathbf{\gamma}$ is fixed, the problem reduces to a standard LASSO formulation, which can be efficiently solved.
For optimizing $\mathbf{\gamma}$, we employ coordinate-wise proximal gradient descent, as described by Parikh and Boyd (2014). The proximal gradient algorithm is a powerful tool for handling non-differentiable convex optimization problems. It decomposes the objective function into two components: a differentiable part and a non-differentiable part, allowing for iterative updates that combine gradient descent with the proximal operator. The proximal operator serves to handle the non-differentiable component, promoting sparsity and effectively reducing the influence of irregularities in the model. By applying the proximal gradient algorithm in our optimization process, we can efficiently minimize the objective function even when the problem involves non-smooth terms, such as LASSO regularization.
@@
$\left(\nabla g(\gamma)\right)_{k}$ $=-\frac{1}{2}\mathbf{r}^{T}\mathbf{\Sigma^{-1}}\left(\mathbf{\left(X_{k}X_{k}^{T}\right)}\odot\mathbf{V}\right)\mathbf{\Sigma^{-1}}\mathbf{r}-\frac{1}{2}q_{k}\|\mathbf{X_{k}^{T}\Sigma^{-1}r}\|^{2}+\frac{1}{2}\operatorname{tr}\left(\left(\mathbf{\left(X_{k}X_{k}^{T}\right)}\odot\mathbf{V}\right)\mathbf{\Sigma}^{-1}\right)$
@@
The proximal algorithm here steps in the direction $\nabla g(\gamma)$ but setting any values that are close to zero equal to zero. For $\sigma^{2}_{0}$, there is no penalty on the parameter. Therefore, we use gradient descent to update $\sigma_{0}^{2}$. To avoid negative values of $\sigma_{0}^{2}$, we use $\tau_{0}=\log(\sigma_{0}^{2})$ for optimization. The gradient of $\tau_{0}$ is given by:
@@
We initialize $\beta$ and $\gamma$ as 0. In each iteration, we first update the covariance matrix $\Sigma$ and transform $\mathbf{X}$ and $\mathbf{Y}$ with the current $\Sigma^{-1/2}$. Then we update $\beta$ by applying LASSO on the transformed data. Then we update $\gamma$ using the proximal algorithm in an elementwise manner. After that, we update $\tau_{0}^{2}$ using gradient descent. The algorithm is summarized in Algorithm 1.
Algorithm 1 Optimization with $\beta$, $\gamma$ and $\tau_{0}$
@@
10:       calculate $\left(\nabla g(\gamma)\right)_{k}$
@@ 3.2 Model selection
  • Cross-validation for $\lambda_{1}$: For each fixed value of $\lambda_{2}$, we use cross-validation to find the best $\lambda_{1}$. We use the cv.glmnet function from the R package glmnet to efficiently perform cross-validation here. We use cross-validation because cv.glmnet allows for quick and easy computation of cross-validation results, making it more convenient and efficient for selecting the optimal parameters.
@@
We choose BIC over other criteria based on empirical findings in Zhang et al. (2024), where pBIC was found to be a more conservative criterion. While a conservative selection method can be beneficial in some cases, detecting shifts in variance is inherently more challenging than detecting shifts in optimal values. A less conservative criterion like BIC is preferable here, as it increases sensitivity to shifts in variance while still penalizing overly complex models. Additionally, BIC offers a computational advantage over pBIC.
@@ 4.1 Comparison of performance
In the main simulation settings, we set $\alpha=1$ and the original variance $\sigma^{2}_{0}=2$. These parameter choices represent a moderate level of phylogenetic signal and baseline trait variance. For each parameter setting, we conducted 50 simulations. To assess the robustness of our conclusions, we also conducted additional simulations with different values of $\alpha$, which are provided in the supplementary materials. The results of these additional simulations are consistent with the main findings presented here. In the following simulations, $\alpha$ is assumed to be unknown. For ShiVa, we obtain $\alpha$ by fitting a null model (without shifts) using the phylolm package.
@@
We compared the performance of different methods based on three key metrics: True Positives (TP), False Positives (FP), predictive log-likelihood, and computational time. Here, True Positives (TP) refer to the number of true shifts correctly detected, while False Positives (FP) refer to the number of false shifts incorrectly detected. However, the positions of shifts are not always identifiable—for example, under a 3-branch tree, any two shifts can lead to an identical model. In such cases, TP and FP may not be reliable metrics, as they would count the true model as including one false positive and missing one true positive, even though the models are effectively equivalent. Figure 3 illustrates the TP and FP results across the various methods under different scenarios. When there is only a shift in the optimal value, ShiVa achieves a well-balanced trade-off between True Positives and False Positives. It effectively detects shifts in the optimal value without generating excessive False Positives. The performance of ShiVa is comparable to other methods that only consider shifts in the optimal value even if only shift in optimal value is present. In contrast, PCMFit produces a much higher number of False Positives, particularly when the signal size is small.
@@
In the case of shifts in variance, ShiVa performs well in detecting variance shifts effectively while maintaining control over False Positives. Although PCMFit shows a higher True Positive rate compared to ShiVa, it comes at the cost of generating significantly more False Positives. Methods like $\ell$1ou and phyloEM, which are designed to detect shifts in optimal value, face challenges when a shift in variance is present. As the magnitude of the variance shift increases, these methods incorrectly interpret the variance signal as multiple shifts in the optimal value. Consequently, their False Positives increase substantially. In contrast, ShiVa successfully distinguishes between variance and mean shifts, effectively controlling the False Positive rate for optimal value shifts, even as the variance shift signal strengthens.
@@
Figure 3: Comparison of True positives and False positives of different methods
In addition to TP and FP, we also evaluated predictive log-likelihood to assess the prediction accuracy of the models estimated by different methods. Regarding identifiability, the OU model is identifiable when no shifts are present. However, previous studies have shown that the exact number and locations of shifts along an edge cannot always be fully determined due to identifiability issues. For example, any two of the three branches connected to a node can yield equivalent models. Even if the selected shifts do not match the true model exactly, choosing a close surrogate shift may be preferable to missing the shift entirely. In such cases, the true positive versus false positive framework might misrepresent performance: a method selecting a surrogate shift is penalized with a false positive, even though it provides a more reasonable approximation than failing to detect the shift. Predictive log-likelihood addresses this issue by evaluating how well the selected models predict new data. For each simulation, we generate 1,000 test datasets for each of the 2,000 simulated datasets, using the same tree and shift configurations. We then compute the mean log-likelihood across the 1,000 test datasets based on the estimated shifts from the training sets. Figure 4 illustrates the differences in predictive log-likelihood between the estimated models and the true model. When shifts occur only in optimal value, ShiVa performs comparably to $\ell$1ou and phyloEM, demonstrating similar predictive precision. However, when shifts in variance are introduced, ShiVa shows a notable improvement, achieving higher predictive log-likelihood than the other methods. In contrast, PCMFit consistently exhibits the lowest predictive log-likelihood across all scenarios, likely due to overfitting.
@@
Average Computation Time Shift Position (mean) Beta Shift Position (variance) Gamma ShiVa l1ou-pBIC l1ou-BIC phyloEM PCMFit 71 1 0 0 95.38 8.09 8.31 68.15 11484.33 71 3 0 0 77.39 8.29 8.57 70.10 13592.56 71 5 0 0 73.86 8.59 8.78 69.92 12227.06 71 -1 0 0 143.39 7.86 8.24 67.81 12890.20 71 -3 0 0 75.70 8.31 8.45 70.02 16215.21 71 -5 0 0 70.33 8.37 8.57 69.35 11962.69 0 0 195 1 70.22 8.08 8.30 68.63 15171.28 0 0 195 1.5 83.31 8.08 8.42 67.72 15680.57 0 0 195 2 77.95 8.04 8.32 70.84 20622.02 0 0 195 3 87.73 8.52 8.81 70.93 16349.04 0 0 195 5 90.05 8.10 8.34 67.70 17358.66 0 0 195 7 97.68 8.18 8.53 67.85 19644.00 0 0 195 -1 101.92 8.04 8.30 67.95 15269.57 0 0 195 -1.5 102.83 7.58 7.74 65.42 15557.74 71 5 195 5 79.32 8.30 8.68 68.18 22836.10 71 3 195 5 129.45 8.55 8.73 68.68 24499.03 71 1 195 5 75.43 8.26 8.92 69.12 18023.39 71 -1 195 5 85.47 7.93 8.24 69.71 19622.12 71 -3 195 5 68.37 7.88 8.51 69.10 24580.25 71 -5 195 5 77.10 7.89 8.21 70.45 23177.43 71 5 195 1 76.73 8.23 8.53 70.06 16574.76 71 5 195 1.5 72.93 8.27 8.59 69.33 13153.55 71 5 195 2 76.58 8.18 8.61 68.71 15742.43 71 5 195 3 90.45 8.17 8.54 69.52 22800.45 71 5 195 7 87.67 8.42 8.67 69.44 23418.19 71 5 195 -1 92.83 8.23 8.45 70.31 16642.62 71 5 195 -1.5 95.15 8.31 8.51 70.79 17517.06

### v3

@@ 2.2 Evolutionary shifts in optimal value and variance
Let ${\gamma}_{b}=\Delta\sigma_{b}^{2}$, $\mathbf{V}_{i,j}=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}(1-e^{-2\alpha t_{a}})$, $\mathbf{U}_{i,j}=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}$, $q_{b}=\frac{e^{-2\alpha T}}{2\alpha}\left(1-e^{2\alpha t_{b}}\right)$; ${\gamma}_{b}$ is the shift in variance on branch $b$, $\mathbf{V}$ is the phylogenetic covariance matrix when $\sigma^{2}=1$ (no shift in variance) with fixed root, $\mathbf{U}$ is the phylogenetic covariance matrix when $\sigma^{2}=1$ (no shift in variance) with stationary distributed root. $\mathbf{U}$ and $\mathbf{V}$ only depend on the phylogeny and $\alpha$. The covariance between species $i$ and $j$ can be expressed as the following equation.
$\mathbf{\Sigma}_{i,j}=\begin{cases}\mathbf{V}_{i,j}\sigma_{0}^{2}+\sum_{b}{\gamma}_{b}\mathbf{X}_{ib}\mathbf{X}_{jb}\mathbf{V}_{i,j}+\sum_{b}{\gamma}_{b}q_{b}\mathbf{X}_{ib}\mathbf{X}_{jb}&\text{OU model with fixed root}\\ \mathbf{U}_{i,j}\sigma_{0}^{2}+\sum_{b}{\gamma}_{b}\mathbf{X}_{ib}\mathbf{X}_{jb}\mathbf{V}_{i,j}+\sum_{b}{\gamma}_{b}q_{b}\mathbf{X}_{ib}\mathbf{X}_{jb}&\text{OU model with random root}\end{cases}$
It consists of 3 terms. The first term is the original covariance without any shift. The last 2 terms show the influence of shifts in variance. In addition to the shift sizes $\gamma_{b}$, the second term is influenced by the phylogentic structure ($\mathbf{V}_{i,j}$), higher original covariance between two species leads to larger change in covariance for a fixed shift size (Fig. 2); the third term is influenced by the start time of branch b ($q_{b}$), earlier shifts lead to larger change in covariance between two species (Fig. 2).
Figure 1: Illustration of how original covariance affects the impact of a variance shift. The green rectangle indicates the location of a variance shift of size $\gamma_{b}$. The asterisks mark the two species whose covariance change is being evaluated. In the left panel, these two species have a higher original covariance (i.e., if no shift occurred), resulting in a larger change in covariance under the same shift, compared to the right panel.
Figure 2: Illustration of how the timing of a variance shift affects covariance change. A shift of size $\gamma_{b}$ is again marked by a green rectangle. The two asterisk-marked species share the same relationship in both panels, but in the left panel, the shift occurs earlier in their shared evolutionary history, leading to a larger change in their covariance than the same shift applied later, as in the right panel.
@@
$=\sigma_{0}^{2}\mathbf{R}+\left(\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma})\mathbf{X^{T}}\right)\odot\mathbf{V}+\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma}\odot\mathbf{q})\mathbf{X^{T}}$ (3)
where $\odot$ denotes elementwise multiplication, and $\operatorname{diag}(\cdot)$ denotes a diagonal matrix with the entries of the vector inside the parentheses on its diagonal. The trait values at tips can be written as:
@@ 3 Methods
$=-\frac{1}{2}(\mathbf{Y}-\beta_{0}-\mathbf{X}\mathbf{\beta})^{T}\left(\sigma_{0}^{2}\mathbf{R}+\left(\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma})\mathbf{X^{T}}\right)\odot\mathbf{V}+\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma}\odot\mathbf{q})\mathbf{X^{T}}\right)^{-1}(\mathbf{Y}-\beta_{0}-\mathbf{X}\mathbf{\beta})$
$\qquad-\frac{1}{2}\log{\det{\left(\sigma_{0}^{2}\mathbf{R}+\left(\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma})\mathbf{X^{T}}\right)\odot\mathbf{V}+\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma}\odot\mathbf{q})\mathbf{X^{T}}\right)}}$
To select the shifts in optimal values ($\beta_{i}\neq 0$) and the shifts in variance ($\gamma_{i}\neq 0$), we use an L1 penalty in the loss function to conduct the feature selection as in LASSO. Therefore, the loss function to be optimized is given by:
@@
$\qquad+\lambda_{1}\|\boldsymbol{\beta}\|_{1}+\lambda_{2}\|\boldsymbol{\gamma}\|_{1}$ (4)
@@ 3.1 Optimization
In this paper, we do not estimate the parameter $\alpha$ jointly with other parameters, but instead treat it as fixed using an ad-hoc estimate obtained from the null model fitted by the phylolm R package. Therefore, in this section, we treat it as fixed during the model fitting process. In the simulations, we demonstrate that the estimation of $\alpha$ does not significantly affect the detection results. The primary objective here is to find the parameters $\boldsymbol{\beta}$, $\boldsymbol{\gamma}$, and $\sigma_{0}^{2}$ that minimize the loss function described in Equation 3. When $\boldsymbol{\gamma}$ is fixed, the problem reduces to a standard LASSO formulation, which can be efficiently solved.
For optimizing $\boldsymbol{\gamma}$, we employ coordinate-wise proximal gradient descent, as described by Parikh and Boyd, (2014). The proximal gradient algorithm is a powerful tool for handling non-differentiable optimization problems. It decomposes the objective function into two components: a differentiable part and a non-differentiable part, allowing for iterative updates that combine gradient descent with the proximal operator. The proximal operator serves to handle the non-differentiable component, promoting sparsity and effectively reducing the influence of irregularities in the model. By applying the proximal gradient algorithm in our optimization process, we can efficiently minimize the objective function even when the problem involves non-smooth terms, such as LASSO regularization.
@@
$\nabla g(\gamma_{k})$ $=-\frac{1}{2}\mathbf{r}^{T}\mathbf{\Sigma^{-1}}\left(\mathbf{\left(X_{k}X_{k}^{T}\right)}\odot\mathbf{V}\right)\mathbf{\Sigma^{-1}}\mathbf{r}-\frac{1}{2}q_{k}\|\mathbf{X_{k}^{T}\Sigma^{-1}r}\|^{2}+\frac{1}{2}\operatorname{tr}\left(\left(\mathbf{\left(X_{k}X_{k}^{T}\right)}\odot\mathbf{V}\right)\mathbf{\Sigma}^{-1}\right)$
@@
The proximal algorithm here steps in the direction $\nabla g(\gamma_{k})$ but setting any values that are close to zero equal to zero. For $\sigma^{2}_{0}$, there is no penalty on the parameter. Therefore, we use gradient descent to update $\sigma_{0}^{2}$. To avoid negative values of $\sigma_{0}^{2}$, we use $\tau_{0}=\log(\sigma_{0}^{2})$ for optimization. The gradient of $\tau_{0}$ is given by:
@@
We initialize $\boldsymbol{\beta}$ and $\boldsymbol{\gamma}$ as $\mathbf{0}$. In each iteration, we first update the covariance matrix $\mathbf{\Sigma}$ and transform $\mathbf{X}$ and $\mathbf{Y}$ with the current $\mathbf{\Sigma}^{-1/2}$. Then we update $\boldsymbol{\beta}$ by applying LASSO on the transformed data. Then we update $\boldsymbol{\gamma}$ using the proximal algorithm in an elementwise manner. After that, we update $\tau_{0}^{2}$ using gradient descent. The algorithm is summarized in Algorithm 1.
Algorithm 1 Optimization with $\boldsymbol{\beta}$, $\boldsymbol{\gamma}$ and $\tau_{0}$
@@
10:         calculate $\nabla g(\gamma_{k})$
@@ 3.2 Model selection
  • Cross-validation for $\lambda_{1}$: For each fixed value of $\lambda_{2}$, we use cross-validation to find the best $\lambda_{1}$. We first de-correlate the data with the estimated covariance matrix. We then use the cv.glmnet function from the R package glmnet to efficiently perform cross-validation here. We use cross-validation because cv.glmnet allows for quick and easy computation of cross-validation results, making it more convenient and efficient for selecting the optimal parameters.
@@
We choose BIC over other criteria based on empirical findings in Zhang et al., (2024), where pBIC was found to be a more conservative criterion. While a conservative selection method can be beneficial in some cases, detecting shifts in variance is inherently more challenging than detecting shifts in optimal values. A less conservative criterion like BIC is preferable here, as it increases sensitivity to shifts in variance while still penalizing overly complex models.
@@ 4.1 Comparison of performance
We generated simulation data under an Ornstein–Uhlenbeck (OU) model with a fixed root at a value $y_{0}=\theta_{0}=0$, using the Anolis lizard phylogeny from Mahler et al., 2013b . This is an ultrametric tree and we rescaled the tree height to 1. The selection strength parameter was set to $\alpha=1$, which corresponds to a phylogenetic half-life of approximately $t_{1/2}=\ln(2)/\alpha\approx 0.693$. The stationary diffusion variance to $\sigma_{0}^{2}=2$. Each parameter setting was repeated across 50 simulation replicates. To assess the robustness of our conclusions, we also performed additional simulations with different values of $\alpha$, which are presented in the supplementary materials and yield results consistent with the main findings. In the following simulations, $\alpha$ is assumed to be unknown. For ShiVa, we estimate $\alpha$ by fitting a null model (without shifts) using the phylolm package.
We considered three simulation scenarios: (1) a single shift in the optimal value, (2) a single shift in the variance, and (3) one shift in the optimal value combined with one shift in the variance.
For the optimal value shift, we varied the shift size across the values $-5$, $-3$, $-1$, $1$, $3$, and $5$. For the variance shift scenarios, we varied the shift magnitude $\gamma_{b}$ across the values $-1.5$, $-1$, $1$, $1.5$, $2$, $3$, $5$, and $7$. In the combined scenario, we first fixed the variance shift at $\gamma_{b}=5$ and varied the optimal value shift as in Scenario (1), and then fixed the optimal value shift at $5$ and varied the variance shift as in Scenario (2). For the setting of $\gamma_{b}$, the largest magnitude considered was 7, which corresponds to a variance that is 3.5 times greater than the original value, given that the baseline $\sigma^{2}$ is 2. In the case study section, we observe examples where the estimated variance shift reaches up to 7 times the original value. Therefore, we believe that the simulated shift magnitudes fall within a realistic and meaningful range.
@@
We compared the performance of different methods using four key metrics: True Positives (TP), False Positives (FP), predictive log-likelihood, and computational time. We count the number of correctly detected true shifts as True Positives (TP), and the number of detected shifts that do not correspond to any true shift as False Positives (FP). The positions of shifts are not always identifiable: for example, under a 3-branch tree, any two shifts can lead to an identical model. In such cases, TP and FP may not be reliable metrics, as they would count the true model as including one false positive and missing one true positive, even though the models are effectively equivalent. However for the simulation studies presented here, the true shifts are identifiable. Figure 3 illustrates the TP and FP results across the various methods under different scenarios. When there is only a shift in the optimal value, ShiVa achieves a well-balanced trade-off between True Positives and False Positives. It effectively detects shifts in the optimal value without generating excessive False Positives. The performance of ShiVa is comparable to other methods that only consider shifts in the optimal value even if only a shift in optimal value is present. In contrast, PCMFit produces a much higher number of False Positives, particularly when the signal size is small.
@@
In the case of shifts in variance, ShiVa performs well in detecting variance shifts effectively while maintaining control over False Positives. Although PCMFit shows a higher True Positive rate compared to ShiVa, it comes at the cost of generating significantly more False Positives. Methods like $\ell$1ou and phyloEM, which are designed to detect shifts in optimal value, face challenges when a shift in variance is present. As the magnitude of the variance shift increases, these methods incorrectly interpret the variance signal as multiple shifts in the optimal value. Consequently, their False Positives increase substantially. In contrast, ShiVa successfully distinguishes between variance and mean shifts, effectively controlling the False Positive rate for optimal value shifts, even as the variance shift signal strengthens. We have more simulations with varying $\alpha$ values in the Supplementary Material and the conclusion remains similar.
Figure 4 shows the detection frequencies of false positive shifts in the optimal value when a true variance shift is present on branch 195 with shift magnitude $\gamma_{b}=7$. The results indicate that the true variance shift on branch 195 is often misinterpreted as shifts in the optimal value on its descendant branches by methods that only consider optimal value shifts, such as $\ell$1ou and phyloEM. PCMFit tends to produce false positive detections more broadly across the tree. In contrast, ShiVa, which models shifts in diffusion variance, is less likely to produce false positive detections of optimal value shifts.
@@
Figure 3: Comparison of True positives and False positives of different methods. Note: In panel (a, left), when $\beta=-5$ or $\beta=5$, all methods correctly detect the shift (True Positives = 1), resulting in overlapping points.
Figure 4: Detection frequencies of false positive shifts in the optimal value under a true variance shift. The pink box highlights the true variance shift on branch 195 (with shift magnitude $\gamma_{b}=7$). Branch 195 refers to the edge label number of the shifted branch.
In addition to TP and FP, we also evaluated predictive log-likelihood to assess the prediction accuracy of the models estimated by different methods. Regarding identifiability, the OU model is identifiable when no shifts are present. However, previous studies have shown that the exact number and locations of shifts along an edge cannot always be fully determined due to identifiability issues. For example, any two of the three branches connected to a node can yield equivalent models. Even if the selected shifts do not match the true model exactly, choosing a close surrogate shift may be preferable to missing the shift entirely. In such cases, the true positive versus false positive framework might misrepresent performance: a method selecting a surrogate shift is penalized with a false positive, even though it provides a more reasonable approximation than failing to detect the shift. To assess how well each method generalizes beyond the training data, we use predictive log-likelihood. Specifically, for each of the 50 training datasets, we estimate model parameters (including the locations and magnitudes of shifts) and use them to compute the expected mean vector and covariance matrix under the fitted model. We then generate 1,000 new test datasets independently using the same phylogeny and true evolutionary process as the training data. For each of these test datasets, we calculate the log-likelihood under the estimated model from the corresponding training set, and take the average to obtain a predictive log-likelihood for that training replicate. This process results in 50 predictive log-likelihood values, from which we report summary statistics (e.g., mean and median). This metric assesses how well the inferred model generalizes to new data drawn from the same evolutionary process, providing insight into the biological relevance of the estimated shifts. Since the log-likelihoods of the true models vary, we compute the predictive log-likelihood difference as the predictive log-likelihood of each estimated model minus the log-likelihood of the corresponding true model. This normalization allows for fair comparison across methods, where a higher score indicates better performance. Figure 5 illustrates the differences in predictive log-likelihood between the estimated models and the true model. When shifts occur only in optimal value, ShiVa performs comparably to $\ell$1ou and phyloEM, demonstrating similar predictive precision. However, when shifts in variance are introduced, ShiVa shows a notable improvement, achieving higher predictive log-likelihood than the other methods. In contrast, PCMFit consistently exhibits the lowest predictive log-likelihood across all scenarios, likely due to overfitting.
@@
Average Computation Time (seconds) Shift Position (mean) Beta Shift Position (variance) Gamma ShiVa l1ou-pBIC l1ou-BIC phyloEM PCMFit 71 1 0 0 95.38 8.09 8.31 68.15 11484.33 71 3 0 0 77.39 8.29 8.57 70.10 13592.56 71 5 0 0 73.86 8.59 8.78 69.92 12227.06 71 -1 0 0 143.39 7.86 8.24 67.81 12890.20 71 -3 0 0 75.70 8.31 8.45 70.02 16215.21 71 -5 0 0 70.33 8.37 8.57 69.35 11962.69 0 0 195 1 70.22 8.08 8.30 68.63 15171.28 0 0 195 1.5 83.31 8.08 8.42 67.72 15680.57 0 0 195 2 77.95 8.04 8.32 70.84 20622.02 0 0 195 3 87.73 8.52 8.81 70.93 16349.04 0 0 195 5 90.05 8.10 8.34 67.70 17358.66 0 0 195 7 97.68 8.18 8.53 67.85 19644.00 0 0 195 -1 101.92 8.04 8.30 67.95 15269.57 0 0 195 -1.5 102.83 7.58 7.74 65.42 15557.74 71 5 195 5 79.32 8.30 8.68 68.18 22836.10 71 3 195 5 129.45 8.55 8.73 68.68 24499.03 71 1 195 5 75.43 8.26 8.92 69.12 18023.39 71 -1 195 5 85.47 7.93 8.24 69.71 19622.12 71 -3 195 5 68.37 7.88 8.51 69.10 24580.25 71 -5 195 5 77.10 7.89 8.21 70.45 23177.43 71 5 195 1 76.73 8.23 8.53 70.06 16574.76 71 5 195 1.5 72.93 8.27 8.59 69.33 13153.55 71 5 195 2 76.58 8.18 8.61 68.71 15742.43 71 5 195 3 90.45 8.17 8.54 69.52 22800.45 71 5 195 7 87.67 8.42 8.67 69.44 23418.19 71 5 195 -1 92.83 8.23 8.45 70.31 16642.62 71 5 195 -1.5 95.15 8.31 8.51 70.79 17517.06